Reyes Corporation applies overhead using a normal costing approach based upon machine-hours. Budgeted factory overhead was $267,840, budgeted machine-hours were 18,600. Actual factory overhead was $288,420, actual machine-hours were 19,150. How much is the over- or underapplied overhead?

Multiple Choice

  • $20,034 underapplied.

  • $12,660 underapplied.

  • $7,374 overapplied.

  • $0.
Answer #1

Predetermined overhead rate=Estimated overhead/Estimated machine hours

=(267840/18600)=14.4 per machine hour

Hence applied overhead=14.4 per machine hour*19150 hours

=275760

Hence since applied overhead is less than actual overhead;

Underapplied overhead=288420-275760

=$12660